The tangent at the point P(θ) on the ellipse x2/4+y2/3 = 1 passes through the point A(2, 1). Show that √3cosθ + sin θ = √3. Find all the solutions of this equation which are in the range 0 ≤ θ < 2π. Hence find the coordinates of the points of contact of the tangents to the ellipse from A. 


1
Expert's answer
2021-08-16T16:48:02-0400
"\dfrac{x^2}{4}+\dfrac{y^2}{3}=1"

"x=2\cos \theta, y=\sqrt{3}\sin \theta"

"y_{x}'=\dfrac{y'_{\theta}}{x'_{\theta}}=-\dfrac{\sqrt{3}\cos \theta}{2\sin \theta}"

"slope=-\dfrac{\sqrt{3}\cos \theta}{2\sin \theta}"

"y-\sqrt{3}\sin \theta=-\dfrac{\sqrt{3}\cos \theta}{2\sin \theta}(x-2\cos \theta)"

"2\sin \theta( y-\sqrt{3}\sin \theta)=-\sqrt{3}\cos \theta(x-2\cos \theta)"

The equation of the tangent line at the point "P(\theta)" to the ellipse is


"x(\sqrt{3}\cos \theta)+y(2\sin \theta)=2\sqrt{3}"

Point "A(2, 1)"


"2(\sqrt{3}\cos \theta)+1(2\sin \theta)=2\sqrt{3}"

"\sqrt{3}\cos \theta+\sin \theta=\sqrt{3}"

Then


"\dfrac{\sqrt{3}}{2}\cos \theta+\dfrac{1}{2}\sin \theta=\dfrac{\sqrt{3}}{2}"

"\sin\dfrac{\pi}{3}\cos \theta+\cos\dfrac{\pi}{3}\sin \theta=\dfrac{\sqrt{3}}{2}, 0\leq\theta\leq2\pi"

"\sin(\dfrac{\pi}{3}+\theta)=\dfrac{\sqrt{3}}{2}"

"\dfrac{\pi}{3}\leq\dfrac{\pi}{3}+\theta\leq\dfrac{7\pi}{3}"

"\dfrac{\pi}{3}+\theta=\dfrac{\pi}{3}=>\theta=0"

Or


"\dfrac{\pi}{3}+\theta=\dfrac{2\pi}{3}=>\theta=\dfrac{\pi}{3}"

Or


"\dfrac{\pi}{3}+\theta=\dfrac{7\pi}{3}=>\theta=2\pi"

"\{0, \dfrac{\pi}{3}, 2\pi\}"